solve the rational equation check extraneous solutions

x2-2x+1/x3+x2-2x=1

-x2+4x/x2-9=4x
Answer #1


- 8 + 1 =) (1) tx - ax a , (y-t) (14) (y-a) 1-) (ча) , ( x-1)(х+а) т Х-) х (х+а) т т (1-1) = (1) (1 ta) – 21 21 +21 -1+= 0Similarly, when & -tih #1 e euthanesu. Hence x = 1 is is extraneous Therefore, x= -1 +iB, x= -1 is are: -*++* -ut 2 29 - 7+ 4